# Hyperfine and Zeeman Optical Pumping and Transverse Laser Cooling of a Thermal Atomic Beam of Dysprosium Using a Single 421 nm Laser

### v2

@@ IV Conclusions and Outlook
Efficient simultaneous optical pumping and laser cooling of atomic dysprosium was achieved with the use of an EOM and a single laser beam. The amplification of the amplitude of the peak of 163Dy in the probe signal after optical pumping is a clear indication that atoms in the hyperfine states were optically pumped into the hyperfine state and the reduced width of the probe spectrum is a clear indicator of laser cooling. Furthermore, a comparison of experimentally obtained amplification in the amplitude of probe spectra with theoretical estimates indicates optimal hyperfine and Zeeman optical pumping with the optical pumping efficiency being constrained by the modulation efficiency of a broadband EOM. The experimentally obtained laser linewidths indicate that the atomic beam has been cooled to the Doppler limit.
The optically pumped dysprosium atoms will be used in the search for parity violation in dysprosium. The previous search for parity violation (PV) [6], the most sensitive parity-violation search so far in terms of the weak-mixing amplitude, has reached into the expected range of the effect but yielded a result consistent with zero. The boost in sensitivity by a factor of afforded by the results of the present work may allow the detection of the long-sought PV signals in Dy.
@@ Acknowledgements.
This research was supported in part by the German Research Foundation (DFG), project ID 390831469: EXC 2118 (PRISMA+ Cluster of Excellence) and project ID BU3035/10-1. JBK is grateful for the support of the National Research Foundation of Korea (NRF-2021R1F1A1060385). JA is grateful for the support of the Humboldt Foundation. D. Antypas and P. Leyser contributed to the project in its early stages. RC is grateful for the continued support and guidance of DB.Appendix A Theory and Simulation
@@
The optical pumping dynamics can be described by the Liouville-von Neumann equations
@@
Simulations 111Simulations are performed using the AtomicDensityMatrix package on Mathematica of Zeeman optical pumping of 164Dy are performed to estimate the time taken to reach a steady state and completely pump the system to the state. Simulations are performed with laser powers used in the experiment and show that the atoms are Zeeman pumped to the state in s. The simulations assume no Doppler broadening and degeneracy of the Zeeman sublevels. Figure 10 shows the time evolution of the population of the ground state Zeeman sublevels at a laser power of 230 mW. Since the experimental interaction times are of the order of 65 s, it is safe to assume that we are fully Zeeman pumping the atomic beam.
@@ Appendix B Estimation of Degree of Optical Pumping
The experimental scheme for measuring the degree of optical pumping relies on measuring the change in the signal strength of the spectral line associated with the optically pumped atoms. The signal strength due to laser-induced fluorescence is proportional to its transition probability which is the square of the transition amplitude. The transition amplitude from the Zeeman sublevel to due to an E1 transition represented by the electric dipole operator () induced by a photon in an atom without nuclear spin is given by
and for atoms with nuclear spin [28, 31],
@@
| 164Dy and 162Dy | 1.94 | - |
| 163Dy and 161Dy | 1.95 | 9 |
It is important to note that Eq. 5 holds only for a single polarization of probe light. Since the Zeeman sublevels are not resolved in the probe spectrum, the amplification of a spectral line in the probe spectrum due to optical pumping can be estimated for even isotopes using
and for the odd isotopes using
where is the population of atoms in a particular Zeeman sublevel and the atoms are initially assumed to be equally distributed among all Zeeman substates. The expected amplification factors in the spectrum of the 598 nm probe inducing transitions with optimal optical pumping are summarised in Table 2.

### v3

@@ IV Conclusions and Outlook
Efficient simultaneous optical pumping and laser cooling of atomic dysprosium was achieved with the use of an EOM and a single laser beam. The amplification of the amplitude of the peak of 163Dy in the probe signal after optical pumping is a clear indication that atoms in the hyperfine states were optically pumped into the hyperfine state and the reduced width of the probe spectrum is a clear indicator of laser cooling. Furthermore, a comparison of experimentally obtained amplification in the amplitude of probe spectra with theoretical estimates indicates optimal hyperfine and Zeeman optical pumping. The experimentally obtained laser linewidths indicate that the atomic beam has been cooled to the Doppler limit. This work has been performed on a atomic beam of Dy but the present technique of using a multitone EOM modulation to pump the hyperfine states of Dy can in principle be used in other configurations such as trapped atoms, molecules and ions where one could expect 100% efficiency in optically pumping the atoms to the polarized states.
The optically pumped dysprosium atoms will be used in the search for parity violation in dysprosium. The previous search for parity violation (PV) [Nguyen1997_PNC_Dy], the most sensitive parity-violation search so far in terms of the weak-mixing amplitude, has reached into the expected range of the effect but yielded a result consistent with zero. The boost in sensitivity, which is proportional to the number of atoms in the state, by a factor of afforded by the results of the present work may allow the detection of the long-sought PV signals in Dy.
@@ Acknowledgements.
This research was supported in part by the German Research Foundation (DFG), project ID 390831469: EXC 2118 (PRISMA+ Cluster of Excellence) and project ID BU3035/10-1. JBK is grateful for the support of the National Research Foundation of Korea (NRF-2021R1F1A1060385). JA is grateful for the support of the Humboldt Foundation. D. Antypas contributed to the project in its early stages. RC is grateful for the continued support and guidance of DB.Appendix A Theory and Simulation
@@
The optical pumping dynamics can be described by the Liouville-von Neumann equations [auzinsh2010opticallypolarisedatoms]
@@
Simulations 111Simulations are performed using the AtomicDensityMatrix package on Mathematica of Zeeman optical pumping of 164Dy are performed to estimate the time taken to reach a steady state and completely pump the system to the state. Simulations are performed with laser powers used in the experiment and show that the atoms are Zeeman pumped to the state in s. The simulations assume no Doppler broadening and degeneracy of the Zeeman sublevels. Fig. 11 shows the time evolution of the population of the ground state Zeeman sublevels at a laser power of 230 mW. Since the experimental interaction times are of the order of 65 s, it is safe to assume that we are fully Zeeman pumping the atomic beam.
@@ Appendix B Estimation of Degree of Optical Pumping
The experimental scheme for measuring the degree of optical pumping relies on measuring the change in the signal strength of the spectral line associated with the optically pumped atoms. The signal strength due to laser-induced fluorescence is proportional to the population in the excited state and the transition probability of the excited state to the ground state.
The excited state population is given by [van_der_Straten_Metcalf_2016],
where is the detuning of the laser from the atomic transition frequency and is the lifetime of the excited state and is the saturation parameter for the transition on resonance. The saturation parameter for a transition is given by [van_der_Straten_Metcalf_2016],
where is the rate of excitation, is the rate of decay from the excited state. The observed fluorescence due to the decay of an excited state would then be proportional to the sum of the transition probabilities through all possible decay paths, and due to the experimental geometry the fluorescence of the decay is maximally observed using the PMT. The observed fluorescence is then given by,
The transition amplitude from the Zeeman sublevel to due to an E1 transition represented by the operator () induced by a photon at a given electric field strength () in an atom without nuclear spin is given by
and for atoms with nuclear spin [auzinsh2010opticallypolarisedatoms, king2008angularmomentumcouplingrabi],
@@
We can deduce from here that the saturation parameter can be represented in terms of Clebsch-Gordan coefficients given by,
where is the laser intensity and is the saturation intensity of the transition. The amplification factor due to complete optical pumping of the hyperfine states to the hyperfine ground state when excited by the 599 nm probe is given by,
where is the population of atoms in a particular excited state Zeeman sublevel. An equal distribution of population is assumed in all the ground state Zeeman sublevels when the atoms are not optically pumped. The saturation intensity of the 599 nm transition is given by, W/mm2 and the intensity of the probe laser used in the experiment is mW/mm2. Fig. 12 shows the variation in the amplification factor with intensity and the amplification factor at intensities used in the experiment are summarized in Table 2.
@@
| 162Dy and 164Dy | 1.2 | - |
| 163Dy and 161Dy | 1.2 | 5.4 |
Appendix C Estimation of number of atoms in the polarized states and amplification of the PNC signal
It is evident from Fig. 5 and 10 that the atoms in the to hyperfine ground states are efficiently optically pumped to the ground state. We know sfrom Fig. 11 that any residual unpumped atoms after the 50 s experimental interaction time will exist in the Zeeman sublevels. It is also a reasonable assumption to make that any residual atoms in the states do not contribute significantly to the amplitude of the probe spectrum due to the reduced density of atoms in these states. With these assumptions, one can estimate the atoms in the state using Eq. 11 and the knowledge of the experimental amplification factor () in the probe spectrum given by,
Assuming an initial unpumped population of one in all the Zeeman sublevels of each hyperfine state, the total number of atoms optically pumped to the state after hyperfine and Zeeman optical pumping with is (). This corresponds to a % efficiency in optical pumping.